Problem 2

A campus researcher wanted to investigate the factors that affect visitor travel time in a complex, multilevel building on campus. Specifically, he wanted to determine whether different building signs (building maps versus wall signage) affect the total amount of time visitors require to reach their destination and whether that time depends on whether the starting location is inside or outside the building. Three subjects were assigned to each of the combinations of signs and starting locations, and travel time in seconds from beginning to destination was recorded. An Excel output of the appropriate analysis is given below:

ANOVA

Source of Variation

SS

df

MS

F

P-value

F crit

Signs

14008.33

14008.33

0.11267

5.317645

Starting Location

12288

2.784395

0.13374

5.317645

Interaction

48

48

0.919506

5.317645

Within

35305.33

4413.167

Total

61649.67

11

A. Referring to the table , the degrees of freedom for the different building signs (factor A) is ____.

  1. 2

  2. 8

  3. 1

  4. 3

B. The mean squares for starting location (Factor B) is ___.

  1. 14,008.3

  2. 48

  3. 12288

  4. 4413.17
Answer #1

F_starting location = MS_starting location / MS_within

2.784395 = MS_starting location / 4413.167

\therefore, MS_starting location (Factor B) = 2.784395 * 4413.167 = 12288

MS_signs (Factor A) = SS_signs / DF_signs

14008.33 = 14008.33 / DF_signs

\therefore DF_signs = 14008.33 / 14008.33 = 1

A. Referring to the table, the degrees of freedom for the different building signs (factor A) is

ANSWER: C. 1

B) The mean squares for starting location (Factor B) is

ANSWER: C. 12288

MS_signs = SS_signs / DF_signs

14008.33 = 14008.33 / DF_signs

\therefore DF_signs = 14008.33 / 14008.33 = 1